OpenStudy (anonymous):

4sin theta= 5cos theta

OpenStudy (anonymous):

\[4 \sin \theta=5 \cos \theta\]

OpenStudy (jdoe0001):

ok.... so.... hmmmm thanks for sharing that

OpenStudy (anonymous):

wait

OpenStudy (jdoe0001):

unless you're meant to find something or solve something

OpenStudy (anonymous):

5/4=sin theta/cos theta

OpenStudy (anonymous):

remember that\[\sin \theta/\cos \theta=\tan \theta\]

OpenStudy (anonymous):

then 5/4= tan theta

OpenStudy (anonymous):

\[\theta=\tan^{-1} (5/4)\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[\theta=51.34^{o}\]

OpenStudy (anonymous):

@jdoe0001 its correct, right?

OpenStudy (jdoe0001):

well... we dunno.. he/she never said what's to be done, proof or getting an angle

OpenStudy (jdoe0001):

or something else

OpenStudy (anonymous):

oh

OpenStudy (kc_kennylau):

as well as the range, theta can be 231.34 degrees also